If the equation describing the motion of a rocket is x=10+5t^2, find its average speed during the first 5.0 seconds of flight.
Answer by stanbon(75887)

If the equation describing the motion of a rocket is h(t)=10+5t^2, find its average speed during the first 5.0 seconds
----------------
Average speed = total distance/total time
-----
Height at t= 0 is 10 ft
Height at t= 5 = 10+5*25 = 135 ft
---
Average speed = (135-10)/(5-0) = 125/5 = 25 ft. per second
======================
Cheers,
Stan H.